OpenStudy (anonymous):

Solve the equation. 64 = 2^(2+x) The "^" indicates that it is an exponent (so 2+x are exponents and are above the 2)

Nnesha (nnesha):

okay 2^what exponent = 64 ??

Nnesha (nnesha):

how many times you should multiply 2 to get 64 ? write 2^exponent at the left side

OpenStudy (er.mohd.amir):

write 64 in exponent form

OpenStudy (er.mohd.amir):

then 64=2 's power some thing then compare powers which give value of x .

OpenStudy (anonymous):

64=2^(2+x) take ln for both sides ln( 64)=ln (2^(2+x) ln(64)=(2+x) ln (2)

OpenStudy (anonymous):

can you solve it? i think so

OpenStudy (anonymous):

If your equation is 64 = (2+x) to the second power then just follow the order of operations. Parenthesis, exponents, multiplication, division, addition, subtraction. when x stands alone you can always put a one in font of it. You can also switch numbers around as long as it doesn't effect the original outcome of the problem. So by doing these things you can make the problem easier to solve by making it (1x+2)^2=64. This should be easier to solve. first undo the addition in the parenthesis by subtracting 2 on both sides of the =. you should get 1x^2=62. next, divide by 1 on both sides of the = to get x^2=62. Then, to finsih the problem and get a final result, you have to figure out what number times itself will equal 62.